OpenStudy (ohmybookness):

a: what is the length of the hypotenuse of triangle ABC? b: find the length of the shorter leg of triangle ABC using two different methods. Show your work for both.

OpenStudy (ohmybookness):

@triciaal this should be easier

OpenStudy (ohmybookness):



jhonyy9 (jhonyy9):

do you know what mean cos35° = ?

OpenStudy (ohmybookness):

is that the correct equation?

jhonyy9 (jhonyy9):

for hypothenuse yes sure

OpenStudy (ohmybookness):

.81

jhonyy9 (jhonyy9):

how you have got it ?

OpenStudy (ohmybookness):

just calculated it

jhonyy9 (jhonyy9):

can you write your calculi here please ?

OpenStudy (ohmybookness):

for method one i got 8.6 so shouldn't my method 2 equal 8.6 too.

jhonyy9 (jhonyy9):

how is this method 2 ?

OpenStudy (ohmybookness):

i have to solve this problem 2 ways.

OpenStudy (ohmybookness):

and show my work

jhonyy9 (jhonyy9):

ok. i understand it - so one way is what i have wrote to you firstly above - yes ?

OpenStudy (ohmybookness):

cos35 degress?

jhonyy9 (jhonyy9):

ok. but first of b. you need to solve a. and this mean that you know the length of hypotenuse - so than you can using for shorter side sin35 = ? and like second method use ?

jhonyy9 (jhonyy9):

cos 35 not can using because cos saying adjacent side divide hypothenuse

jhonyy9 (jhonyy9):

and you need to calculi the length of shorter side - yes ?

jhonyy9 (jhonyy9):

tan 35 equal with ? do you know what say tan of an angle ?

OpenStudy (ohmybookness):

hmm

OpenStudy (ohmybookness):

.7?

OpenStudy (ohmybookness):

sorry. I'm not good at math

OpenStudy (triciaal):

It might be helpful to note that angles in a triangle att to 180. So the other angle is 55 Tan 55 = 15/BC solve for BC the shorter side

OpenStudy (ohmybookness):

wait where'd you get the 55 :(

OpenStudy (triciaal):

Sin Rule a/Sin A = b/Sin B = c/Sin C

OpenStudy (triciaal):

I don't know how to draw on the iPad

OpenStudy (ohmybookness):

oh no! :(

OpenStudy (ohmybookness):

okay so far I wrote down Tan55=15/bc. is BC, B times C?

OpenStudy (triciaal):

Typo above Angles in a triangle ADD up to 180

OpenStudy (triciaal):

The line from B to C

OpenStudy (ohmybookness):

oh ok

OpenStudy (ohmybookness):

so is the equation 1.42=15/BC ?

OpenStudy (ohmybookness):

someone from open study showed me another method because my school wants me to solve it using two methods, and as an answer I got 8.6. would that be correct?

OpenStudy (triciaal):

Need to get to my computer so I can draw. Will check if you are still "open"

OpenStudy (ohmybookness):

got it

OpenStudy (triciaal):

What did you get when you solved the equation for the line BC ?

OpenStudy (triciaal):

Regardless of the method you should get the same length

OpenStudy (ohmybookness):

I got 1.42=15/BC

OpenStudy (triciaal):

Think of BC as the unknown single variable and solve for it

OpenStudy (ohmybookness):

I got 10.71

OpenStudy (triciaal):

Do you see that BC = 15 divided by 1.42 based on your equation?

OpenStudy (ohmybookness):

I got 10.71